# Segmentation-free Goodness of Pronunciation

### v2

@@ Abstract
Mispronunciation detection and diagnosis (MDD) is a significant part in modern computer aided language learning (CALL) systems. Within MDD, phoneme-level pronunciation assessment is key to helping L2 learners improve their pronunciation. However, most systems are based on a form of goodness of pronunciation (GOP) which requires pre-segmentation of speech into phonetic units. This limits the accuracy of these methods and the possibility to use modern CTC-based acoustic models for their evaluation. In this study, we first propose self-alignment GOP (GOP-SA) that enables the use of CTC-trained ASR models for MDD. Next, we define a more general alignment-free method that takes all possible alignments of the target phoneme into account (GOP-AF). We give a theoretical account of our definition of GOP-AF, an implementation that solves potential numerical issues as well as a proper normalization which makes the method applicable with acoustic models with different peakiness over time. We provide extensive experimental results on the CMU Kids and Speechocean762 datasets comparing the different definitions of our methods, estimating the dependency of GOP-AF on the peakiness of the acoustic models and on the amount of context around the target phoneme. Finally, we compare our methods with recent studies over the Speechocean762 data showing that the feature vectors derived from the proposed method achieve state-of-the-art results on phoneme-level pronunciation assessment.
Abstract
Mispronunciation detection and diagnosis (MDD) is a significant part in modern computer aided language learning (CALL) systems. Within MDD, phoneme-level pronunciation assessment is key to helping L2 learners improve their pronunciation. However, most systems are based on a form of goodness of pronunciation (GOP) which requires pre-segmentation of speech into phonetic units. This limits the accuracy of these methods and the possibility to use modern CTC-based acoustic models for their evaluation. In this study, we first propose self-alignment GOP (GOP-SA) that enables the use of CTC-trained ASR models for MDD. Next, we define a more general alignment-free method that takes all possible alignments of the target phoneme into account (GOP-AF). We give a theoretical account of our definition of GOP-AF, an implementation that solves potential numerical issues as well as a proper normalization which makes the method applicable with acoustic models with different peakiness over time. We provide extensive experimental results on the CMU Kids and Speechocean762 datasets comparing the different definitions of our methods, estimating the dependency of GOP-AF on the peakiness of the acoustic models and on the amount of context around the target phoneme. Finally, we compare our methods with recent studies over the Speechocean762 data showing that the feature vectors derived from the proposed method achieve state-of-the-art results on phoneme-level pronunciation assessment.
Index Terms:
Computer-aided pronunciation training, mispronunciation detection and diagnosis, speech assessment, goodness of pronunciation, CTC, child speech, L2.
Index Terms:
Computer-aided pronunciation training, mispronunciation detection and diagnosis, speech assessment, goodness of pronunciation, CTC, child speech, L2.
@@ I Introduction
Witt et al. [witt2000phone] proposed a widely used method for MDD at the phoneme level that is based on a measure called goodness of pronunciation (GOP). The advantage of this method is that it relies on acoustic models exclusively trained for automatic speech recognition (ASR). The ASR models are used to score new utterances, and only a small amount of MDD annotations are required to optimize a detection threshold that separates acceptable pronunciations from mispronunciations. ASR models can also be used to transcribe speech at the word, character, or phoneme level and detect mispronunciations by comparing the resulting output sequence with the canonical pronunciation or orthographic transcription, respectively [cnn-rnn-ctc, fu2021textdependent, textdep2, hybrid, getman2022wav2vec2].
Another approach is to train end-to-end models for MDD. However, training such models from scratch would require large amounts of MDD annotated data. A solution is to use foundation models that are either trained on large amounts of unlabeled speech data, or fined tuned for ASR. These models can be further fine-tuned with small amounts of MDD annotated data for the MDD task. For example, Xu et al.[xu2021explore] follow this strategy starting from a Wav2Vec2.0 model [w2v2] whereas, the authors of [ssl1, ssl2] start from a HuBERT model [hubert]. Liu et.al. [Liu2023ZeroShotAP] attempt to perform MDD at utterance-level based on hidden representations directly derived from the foundation models without further fine-tuning.
In this paper, our goal is to combine the advantages of the GOP method, the foundation models, and end-to-end ASR training based on CTC loss [CTC]. This raises a number of challenges due to the fact that GOP requires segmentation of speech into phonetic units. This is typically obtained by forced-alignment of the canonical pronunciation with the spoken utterance. If the pronunciation is correct, the obtained phonetic boundaries may vary due to co-articulation effects. In case of mispronunciations, the segmentation may be even less reliable. Finally, CTC trained models tend to give activations that are not necessarily aligned with acoustic speech segmentation. In [cao24b_interspeech] we introduced a framework for combining CTC trained models and GOP. In particular, we introduced 1) a self-aligned GOP method that uses the same activations from CTC trained models for alignment and GOP evaluation, and 2) an alignment free GOP method that can assess a specific speech segment without committing to a particular segmentation.
In this paper we enhance these method by making the following novel contributions:
  • We define proper segment length normalization for our alignment-free GOP definition without committing to a particular segmentation.
  • We provide a theoretical derivation of the alignment-free GOP method that exposes the assumptions required for its definition.
@@
  • We provide extensive experimental results that compare the different methods we propose to the state-of-the-art on binary and three levels MDD tasks.
  • We provide an analysis of the effect of peakiness of the model on the alignment-free GOP method;
  • We provide an analysis of the effect of context length (utterance length) on the alignment-free GOP method;
  • We will make all code available upon publication.
The paper is organized as follows: in Section II we provide the background and review the related work that are necessary to understand our methods. In Section III we give our GOP definitions and theoretical derivations showing the assumptions that are implicit in them. Sections IV and V report on our experimental settings and empirical results. Finally, Section VI concludes the paper.
Figure 1: Standard GOP methods use forced-alignment to segment speech for analysis.
@@ II Background and Related Work
In this section we give background information and review the related work that are necessary to understand the proposed method. We will focus on phoneme-level pronunciation assessment that is the goal of this paper. We first need to distinguish between continuous measures for pronunciation assessment and the specific task of mis-pronunciation detection and diagnosis (MDD). In the first case, the goal is to introduce a metric that can indicate how far a specific pronunciation of a phoneme is from the acceptable variability in the language. An example of this kind of measure is goodness of pronunciation (GOP). The MDD task, on the other hand, is to provide a binary or, sometimes ternary, decision on whether the pronunciation is correct or not. This task can be performed 1) on a single GOP-like score by defining thresholds for the different output classes, 2) by more advanced classifiers based on multidimensional engineered features, or 3) by end-to-end systems that take raw speech as input.
@@ II-A The definition of GOP
Goodness of pronunciation (GOP) was initially proposed as a measure of how closely the pronunciation of a specific phoneme is to its expected canonical pronunciation [witt2000phone]. This is a segmental measure that requires time-aligned phonetic annotations (see e.g. Figure 1). Witt’s original definition of GOP [witt2000phone] corresponds to the log posterior of the canonical phoneme $l_{i}$ given the sequence of observations $O_{t_{1}}^{t_{2}}=\{o_{t_{1}},\dots,o_{t_{2}}\}$ in the segment under test, normalized by the sequence length:
@@
The estimation of the posteriors was originally implemented with Hidden Markov Models and Gaussian Mixture Models (HMM-GMM) and by using Bayes rule:
@@
The phoneme boundaries $t_{1}$ and $t_{2}$ could be obtained by human annotations. However, this is not practical: During model training, annotating large amounts of data would be too costly. More importantly, this would make the methods not suitable for providing immediate and automated feedback to the students. The segmentation step is, therefore, commonly automated by forced-alignment using the canonical pronunciation $L_{\text{C}}=\{l_{1},\dots,l_{|L_{\text{C}}|}\}$ and a phonetic model trained for ASR. This model does not need to be the same used for the GOP evaluation, and it is typically a context-independent HMM-GMM model.
@@
Though the success of Witt’s GOP, many later works argue that the phone-loop approximation is not reliable in estimating the denominator of Eq. 2. For example, Lattice-base-GOP [Song2010LatticebasedGI] includes contributions from N-best hypotheses to avoid under-estimating the value of the denominator. In [luo2009analysis], the authors show performance improvements that can be obtained when the phone-loop is evaluated multiple times over the target segment for each phoneme rather than once along the whole utterance. This implementation follows more closely the original definition of the denominator in Eq.2.
@@
Phonological rules from the learner’s first language (L1) can also be included in the GOP methods, where the acoustic models are trained with target language (L2) only, to achieve better accuracy [ERN, DUDY201862].
@@
The models used for estimating the GOP values are trained for ASR and therefore do not need any data that are annotated specifically for the MDD task. However, in order to detect mispronunciations, a small amount of human-annotated data is required to estimate the optimal threshold that separates good pronunciations from bad ones. These thresholds are usually phoneme-dependent but even finer-grained definitions (for example speaker-dependent) may improve accuracy (e.g. [kanters09_slate]).
@@
Using hybrid HMM-Deep Neural Networks (DNN) acoustic models has been shown to improve MDD performance [hu13_interspeech]. In this case, the posterior defined in Eq. 1 can be evaluated directly from the posterior probabilities $p(l_{i}|o_{t})$ from the softmax output of the last DNN layer for the acoustic frames in the target phoneme. The corresponding GOP definition is then the average of the log posterior probabilities:
[…]
@@
The DNN is often trained frame-wise with cross entropy (CE) loss. Similarly to the HMM-GMM based GOP, the segment boundaries $t_{1}$ and $t_{2}$ are obtained from an external aligner that is also used to provide the frame-wise labels for training the DNN. We call this method external alignment (EA) in the following. In [improved_DNN], the authors show that the performance of DNN-based GOP can be further improved by merging the probabilities of fine-grained sub-phoneme symbols (senones) to phonemes. [^1] To compare with CTC-based methods (GOP-CTC-AF and GOP-CTC-SA) where GOP is also operated at phoneme level, our later discussion and experiments are all based on models trained directly at phoneme level.
@@
In Eq. 3 each frame is evaluated independently using the output of the DNN, neglecting the transition probabilities in the HMM. In [HMM_tran_GOP], the author argues that these probabilities can contribute with contextual information and including them in the GOP evaluation can improve performance. In [context-aware-GOP], transition and duration factors are introduced to increase context awareness in GOP.
@@ II-B The alignment issues for GOP
[…]
Even assuming the existence of well defined phonetic boundaries, alignment errors may occur in case of mispronunciations because the models used for segmentation are usually trained on correctly pronounced speech. Furthermore, as studied in [the_impact_of_FA, perform_FA, L2-GOP], forced-alignment may be unreliable due to speaker variability: age, accent, dialect, health condition, to name a few. The uncertainty over phonetic boundaries has a strong impact on traditional GOP definitions which consider the boundaries as deterministic.
@@
In [GPP] the authors propose to use the general posterior probability (GPP) to mitigate alignment issues by allowing evaluation over any segments that overlap with the target segment. Zhang et al. [textdep3] propose a pure end-to-end method that uses a sequence-to-sequence model without having to segment the speech. However, this method requires a large number of human annotations for MDD and an additional step that compares the canonical phoneme sequence with the human-annotated sequence before training the model.
In this work, we propose two methods to relax the dependency of GOP on the accuracy of segmentation. The first has the goal of reducing the effect of misalignment between the acoustic model activations and the speech segments, the second is completely alignment-free. The latter also allows for insertion and deletion errors that were mentioned in Section II-A.
Figure 2: Illustration of the issues with standard GOP methods for models trained with CE loss (left) and CTC loss (right). Each plot corresponds to a different kind of mispronunciation and is divided in three parts: “sp” shows that segments that were actually spoken; “fa” shows the forced alignment based on the canoncial pronunciation; “ac” shows the activations from the model and for the canonical segments. Finally, for the CTC models, the red dashed lines show the alternative forced alignment from the GOP-CTC-align method.
@@ II-C End-to-end ASR models, CTC and peaky behavior
End-to-end ASR models were initially introduced to map speech to output symbols, typically characters [^2] The methods for MDD described in this paper use phonemes as output symbols. However, all the arguments in this section apply regardless of the nature of the output symbols., directly. In general, the sequence of output symbols $L=\{l_{1},\dots,l_{|L|}\}$ has a different rate that the input speech feature vectors $O_{1}^{T}=\{o_{1},\dots,o_{T}\}$. To overcome this problem in training, the CTC loss [CTC] was introduced. This makes use of an additional blank ($\phi$) output symbol. Given a speech sequence $O_{1}^{T}$ of length $T$, we define $V$ as the set of all output symbols including $\phi$. Then any vector $u=\{u_{1},\dots,u_{T}\}\in U=V^{T}$ is an alignment path between the input sequence and the output symbols. The probability of a path $u$ under the assumption of conditional independence is $p(u|O_{1}^{T})=\prod_{t=1}^{T}p(u_{t}|o_{t})$. As a consequence $\sum_{u\in U}p(u|O_{1}^{T})=1$ representing all possible paths given the model and the speech. During training, for a given target output sequence $L$ and the speech sequence $O_{1}^{T}$, the model learns to minimize the following loss:
@@
where each $\pi$ is any of the paths that can be mapped to $L$ using the many to one function: $\mathcal{B}:U\rightarrow L$ by removing all the repeated symbols, first and then the blank symbols, e.g., $\mathcal{B}(a{\phi}b{\phi}b{\phi}cc{\phi})$ $=$ $abbc$.
Since the training of the CTC does not require frame-level alignment, the timing information tends to be ignored by the model. A well-known phenomenon of CTC-trained model is the “peaky” behavior of model output, as illustrated in the right column of Figure 2. There are two dimensions of peakiness: “peaky over time” (POT) and “peaky over state” (POS). POT behavior corresponds to the fact that the blank symbol is activated for most of the time steps, whilst the non-blank symbols that correspond to the target label sequence only become activated for a few time steps. Forced-alignment (Viterbi search) performed with CTC-trained models can result in distorted alignment between input speech and output symbols. On the other hand, POS behavior is observed because the model activations at each frame are always close to 1.0 for one output and 0.0 for the all the others.
For these reasons, CTC is not feasible for applications where the time alignment is essential, such as voice activity detection, speech segmentation, or conventional GOP evaluation for MDD. An indication of this is the lack of literature in these areas that use CTC-based models. However, several works tried to mitigate the peaky behavior of CTC-based models, e.g. [why_peaky, CTC-ME, huang2024peakyaccuratectcforced].

### v3

@@ Abstract
Mispronunciation detection and diagnosis (MDD) is a significant part in modern computer-aided language learning (CALL) systems. Most systems implementing phoneme-level MDD through goodness of pronunciation (GOP), however, rely on pre-segmentation of speech into phonetic units. This limits the accuracy of these methods and the possibility to use modern CTC-based acoustic models for their evaluation. In this study, we first propose self-alignment GOP (GOP-SA) that enables the use of CTC-trained ASR models for MDD. Next, we define a more general segmentation-free method that takes all possible segmentations of the canonical transcription into account (GOP-SF). We give a theoretical account of our definition of GOP-SF, an implementation that solves potential numerical issues as well as a proper normalization which allows the use of acoustic models with different peakiness over time. We provide extensive experimental results on the CMU Kids and speechocean762 datasets comparing the different definitions of our methods, estimating the dependency of GOP-SF on the peakiness of the acoustic models and on the amount of context around the target phoneme. Finally, we compare our methods with recent studies over the speechocean762 data showing that the feature vectors derived from the proposed method achieve state-of-the-art results on phoneme-level pronunciation assessment.
@@ I Introduction
Witt et al. [37] proposed a widely used method for MDD at the phoneme level that is based on a measure called goodness of pronunciation (GOP). The advantage of this method is that it relies on acoustic models exclusively trained for automatic speech recognition (ASR). The ASR models are used to score new utterances, and only a small amount of MDD annotations are required to optimize a detection threshold that separates acceptable pronunciations from mispronunciations. ASR models can also be used to transcribe speech at the word, character, or phoneme level and detect mispronunciations by comparing the resulting output sequence with the canonical pronunciation or orthographic transcription, respectively [24, 11, 10, 41, 12].
Another approach is to train end-to-end models for MDD. However, training such models from scratch would require large amounts of MDD annotated data. A solution is to use foundation models that are either trained on large amounts of unlabeled speech data, or fine-tuned for ASR. These models can be further fine-tuned with small amounts of MDD annotated data for the MDD task. For example, Xu et al. [38] follow this strategy starting from a Wav2Vec2.0 model [1] whereas, the authors of [23, 31] start from a HuBERT model [17]. Liu et al. [25] attempt to perform MDD at utterance-level based on hidden representations directly derived from the foundation models without further fine-tuning.
In this paper, our goal is to combine the advantages of the GOP method, the foundation models, and end-to-end ASR training based on CTC loss [14]. This raises a number of challenges due to the fact that GOP requires segmentation of speech into phonetic units. This is typically obtained by forced alignment of the canonical pronunciation with the spoken utterance. If the pronunciation is correct, the obtained phonetic boundaries may vary due to coarticulation effects. In case of mispronunciations, the segmentation may be even less reliable. Finally, CTC trained models tend to give activations that are not necessarily aligned with acoustic speech segmentation. In [4] we introduced a framework for combining CTC trained models and GOP. In particular, we introduced 1) a self-aligned GOP method that uses the same activations from CTC trained models for alignment and GOP evaluation, and 2) a segmentation-free GOP method that can assess a specific speech segment without committing to a particular segmentation.
In this paper, we enhance these methods by making the following novel contributions:
  • We define proper segment length normalization for our segmentation-free GOP definition without committing to a particular segmentation.
  • We provide a theoretical derivation of the segmentation-free GOP method that exposes the assumptions required for its definition.
@@
  • We provide extensive experimental results that compare the different methods we propose to the state-of-the-art on binary and ternary MDD tasks.
  • We provide an analysis of the effect of peakiness of the model on the segmentation-free GOP method.
  • We provide an analysis of the effect of context length (utterance length) on the segmentation-free GOP method.
  • We make all code available. [^1] https://github.com/frank613/CTC-based-GOP
The paper is organized as follows: in Section II we provide the background and review the related work that is necessary to understand our methods. In Section III we give our GOP definitions and theoretical derivations showing the assumptions that are implicit in them. Sections IV and V report on our experimental settings and empirical results. Finally, Section VI concludes the paper.
Figure 1: Standard GOP methods use forced alignment to segment speech for analysis.
@@ II Background and Related Work
In this section we give background information and review the related work that are necessary to understand the proposed method. We will focus on phoneme-level pronunciation assessment that is the goal of this paper. We first need to distinguish between continuous measures for pronunciation assessment and the specific task of MDD. In the first case, the goal is to introduce a metric that can indicate how far a specific pronunciation of a phoneme is from the acceptable variability in the language. An example of this kind of measure is GOP. The MDD task, on the other hand, is to provide a binary or, sometimes ternary, decision on whether the pronunciation is correct or not. This task can be performed 1) on a single GOP-like score by defining thresholds for the different output classes, 2) by more advanced classifiers based on multidimensional engineered features, or 3) by end-to-end systems that take raw speech as input.
@@ II-A The definition of GOP
GOP was initially proposed as a measure of how closely the pronunciation of a specific phoneme matches its expected canonical pronunciation [37]. This is a segmental measure that requires time-aligned phonetic annotations (see, e.g., Figure 1). Witt's original definition of GOP [37] corresponds to the log posterior of the canonical phoneme $l_{i}$ given the sequence of observations $O_{t_{1}}^{t_{2}}=\{o_{t_{1}},\dots,o_{t_{2}}\}$ in the segment under test, normalized by the sequence length:
@@
The estimation of the posteriors was originally implemented using Hidden Markov Models and Gaussian Mixture Models (HMM-GMM) and by using Bayes rule:
@@
The phoneme boundaries $t_{1}$ and $t_{2}$ could be obtained by human annotations. However, this is not practical: During model training, annotating large amounts of data would be too costly. More importantly, this would make the methods not suitable for providing immediate and automated feedback to the students. The segmentation step is, therefore, commonly automated by forced alignment using the canonical pronunciation $L_{\text{C}}=\{l_{1},\dots,l_{|L_{\text{C}}|}\}$ and a phonetic model trained for ASR. This model does not need to be the same as that used for the GOP evaluation, and it is typically a context-independent HMM-GMM model.
@@
Although Witt's GOP has been successful, many later works argue that the phone-loop approximation is not reliable in estimating the denominator of Eq. 2. For example, Lattice-based GOP [33] includes contributions from N-best hypotheses to avoid underestimating the value of the denominator. In [27], the authors show performance improvements that can be obtained when the phone-loop is evaluated multiple times over the target segment for each phoneme rather than once along the whole utterance. This implementation follows more closely the original definition of the denominator in Eq.2.
@@
Phonological rules from the learner's first language (L1) can also be included in the GOP methods, where the acoustic models are trained with target language (L2) only, to achieve better accuracy [16, 8].
@@
The models used for estimating the GOP values are trained for ASR and therefore do not need any data that are annotated specifically for the MDD task. However, in order to detect mispronunciations, a small amount of human-annotated data is required to estimate the optimal threshold that separates good pronunciations from bad ones. These thresholds are usually phoneme-dependent but even finer-grained definitions (for example speaker-dependent) may improve accuracy (e.g. [22]).
@@
Using hybrid HMM-Deep Neural Networks (DNN) acoustic models has been shown to improve MDD performance [18]. In this case, the posterior defined in Eq. 1 can be evaluated directly from the posterior probabilities $p(l_{i}|o_{t})$ from the softmax output of the last DNN layer for the acoustic frames in the target phoneme. The corresponding GOP definition is then the average of the log posterior probabilities:
$\text{GOP-Avg}(l_{i})=\frac{1}{t_{2}-t_{1}}\sum_{t=t_{1}}^{t_{2}}\log\left({\mathit{p}(l_{i}|o_{t})}\right).$ (3)
@@
The DNN is often trained frame-wise with cross entropy (CE) loss. Similarly to the HMM-GMM based GOP, the segment boundaries $t_{1}$ and $t_{2}$ are obtained from an external aligner that is also used to provide the frame-wise labels for training the DNN. We call this method external alignment (EA) in the following. In [20], the authors show that the performance of DNN-based GOP can be further improved by merging the probabilities of fine-grained sub-phoneme symbols (senones) to phonemes. [^2] To compare with CTC-based methods (GOP-CTC-SF and GOP-CTC-SA) where GOP is also operated at phoneme level, our later discussion and experiments are all based on models trained directly at phoneme level.
@@
In Eq. 3 each frame is evaluated independently using the output of the DNN, neglecting the transition probabilities in the HMM. In [34], the author argues that these probabilities can contribute with contextual information and including them in the GOP evaluation can improve performance. In [32], transition and duration factors are introduced to increase context awareness in GOP.
@@ II-B The alignment issues for GOP
There are several issues with traditional definitions of GOP related to speech segmentation. The first is the ability to perform a perfect alignment of phonetic segments to the recorded speech. Phonetic segmentation is an ill-posed problem because it is based on the assumption that speech is produced as a sequence of discrete events. However, coarticulation effects in speech production question the existence of explicit phonetic boundaries. This explains the disagreement on segmentation even between trained phoneticians. An example is shown in Figure 1, where it is difficult to tell the exact beginning and end of the sonorant [w] as it happens in the transition of two vowel sounds [ı] and [].
Even assuming the existence of well-defined phonetic boundaries, alignment errors may occur in case of mispronunciations because the models used for segmentation are usually trained on correctly pronounced speech. Furthermore, as studied in [2, 28, 19], forced alignment may be unreliable due to speaker variability: age, accent, dialect, health condition, to name a few. The uncertainty over phonetic boundaries has a strong impact on traditional GOP definitions which consider the boundaries as deterministic.
@@
In [35] the authors propose to use the general posterior probability (GPP) to mitigate alignment issues by allowing evaluation over any segments that overlap with the target segment. Zhang et al. [42] propose a pure end-to-end method that uses a sequence-to-sequence model without having to segment the speech. However, this method requires a large number of human annotations for MDD and an additional step that compares the canonical phoneme sequence with the human-annotated sequence before training the model.
In this work, we propose two methods to relax the dependency of GOP on the accuracy of segmentation. The first has the goal of reducing the effect of misalignment between the acoustic model activations and the speech segments, the second is completely segmentation-free. The latter also allows for insertion and deletion errors that were mentioned in Section II-A.
Figure 2: Illustration of the issues with standard GOP methods for models trained with CE loss (left) and CTC loss (right). Each plot corresponds to a different kind of mispronunciation and is divided in three parts: ``sp'' shows that segments that were actually spoken; ``fa'' shows the forced alignment based on the canonical pronunciation; ``ac'' shows the activations from the model and for the canonical segments. Finally, for the CTC models, the red dashed lines show the alternative forced alignment from the GOP-CTC-align method.
@@ II-C End-to-end ASR models, CTC and peaky behavior
End-to-end ASR models were initially introduced to map speech to output symbols, typically characters [^3] The methods for MDD described in this paper use phonemes as output symbols. However, all the arguments in this section apply regardless of the nature of the output symbols., directly. In general, the sequence of output symbols $L=\{l_{1},\dots,l_{|L|}\}$ has a different rate that the input speech feature vectors $O_{1}^{T}=\{o_{1},\dots,o_{T}\}$. To overcome this problem in training, the CTC loss [14] was introduced. This makes use of an additional blank ($\phi$) output symbol. Given a speech sequence $O_{1}^{T}$ of length $T$, we define $V$ as the set of all output symbols including $\phi$. Then any vector $u=\{u_{1},\dots,u_{T}\}\in U=V^{T}$ is an alignment path between the input sequence and the output symbols. The probability of a path $u$ under the assumption of conditional independence is $p(u|O_{1}^{T})=\prod_{t=1}^{T}p(u_{t}|o_{t})$. As a consequence $\sum_{u\in U}p(u|O_{1}^{T})=1$ representing all possible paths given the model and the speech. During training, for a given target output sequence $L$ and the speech sequence $O_{1}^{T}$, the model learns to minimize the following loss:
@@
where $\pi$ spans over the paths that can be mapped to $L$ using the many-to-one function: $\mathcal{B}:U\rightarrow L$ by removing all repeated symbols and followed by removal of the blank symbols, e.g., $\mathcal{B}(a{\phi}b{\phi}b{\phi}cc{\phi})$ == abbcabbc.
Since training the CTC does not require frame-level alignment, the timing information tends to be ignored by the model. A well-known phenomenon of CTC-trained model is the ``peaky'' behavior of model output, as illustrated in the right column of Figure 2. There are two dimensions of peakiness: ``peaky over time'' (POT) and ``peaky over state'' (POS). POT behavior corresponds to the fact that the blank symbol is activated for most of the time steps, whilst the non-blank symbols that correspond to the target label sequence only become activated for a few time steps. Forced alignment (Viterbi search) performed with CTC-trained models can result in distorted alignment between input speech and output symbols. On the other hand, POS behavior is observed because the model activations at each frame are always close to 1.0 for one output and 0.0 for all the others.
For these reasons, CTC is not feasible for applications where the time alignment is essential, such as voice activity detection, speech segmentation, or conventional GOP evaluation for MDD. An indication of this is the lack of literature in these areas that use CTC-based models. However, several works tried to mitigate the peaky behavior of CTC-based models for ASR, e.g. [39, 26, 21].